OpenStudy (anonymous):

1. Imagine a local bank representative is assisting you with establishing an account. The account you two are discussing has an APR of 6.5%. Determine the APY with quarterly compounding and with monthly compounding. How does changing the compounding period affect the annual yield?

OpenStudy (anonymous):

quarterly =1/4 monthly =12

OpenStudy (anonymous):

A=P(1+r/n)^(n*t)

OpenStudy (dumbcow):

use that formula above assume we are depositing 1 for 1 year P=1 t=1 r=.065 n=4 or 12 find A in both cases APY = (A-1)/1 * 100%

OpenStudy (anonymous):

does A=3.08, how do you find y

OpenStudy (dumbcow):

i think y or yield is A as we've defined it A should be really close to 1

OpenStudy (anonymous):

I use the formula A=1(1+.065/4) (4*1)

OpenStudy (dumbcow):

oh not times 4 , raised to the 4th power

OpenStudy (dumbcow):

^ means exponent

OpenStudy (anonymous):

I got .0050252592 for A

OpenStudy (dumbcow):

how did you get that? .065/4 =.01625 add 1 = 1.01625 raise to 4th power = 1.0666

OpenStudy (anonymous):

so A =1.0666 APY=.0666

OpenStudy (dumbcow):

times 100%

OpenStudy (anonymous):

i did the answer didnt change